# Bayesian Inference General Procedures for A Single-subject Test Study

### v4

@@ Acknowledgments
This KTP project (ID:10053865, KTP13481) received financial support from UKRI through Innovate UK and Innovision IP Ltd. This work was supported by the High-Performance Computing Cluster at the University of Kent.
@@ Appendix A Appendix
where $r(z,\nu)=\sqrt{(\nu+1)/(\nu+z^{2})}$. Given any $\nu>0$, let $\boldsymbol{\pi}(\alpha|\nu)$ represent the Jeffery’s prior for $\alpha$, we have
@@
By definition of Jeffery’s prior, we have $\boldsymbol{\pi}(\alpha|\nu)\propto\sqrt{I_{\alpha\alpha}}$, where $I_{\alpha\alpha}$ is the Fisher information of $\alpha$ in the skewed Student $t$ density function $f(z|\alpha,\nu)$. According to Branco et al. (2013, Proposition 2), the Fisher information of $\alpha$ in $f(z|\alpha,\nu)$ is given by
@@
Other results in Section 3.2

### v5

@@ 5 Conclusion
Funding
This research received financial support jointly from UKRI through Innovate UK and Innovision IP Ltd. (Grant No: 10053865).
@@ Acknowledgments
This work was supported by the High-Performance Computing Cluster at the University of Kent.
@@ Appendix A Appendix
where $r(z,\nu)=\sqrt{(\nu+1)/(\nu+z^{2})}$. Given any $\nu>0$, let $\boldsymbol{\pi}(\alpha|\nu)$ represent the Jeffreys prior for $\alpha$, we have
@@
By definition of Jeffreys prior, we have $\boldsymbol{\pi}(\alpha|\nu)\propto\sqrt{I_{\alpha\alpha}}$, where $I_{\alpha\alpha}$ is the Fisher information of $\alpha$ in the skewed Student $t$ density function $f(z|\alpha,\nu)$. According to Branco et al. (2013, Proposition 2), the Fisher information of $\alpha$ in $f(z|\alpha,\nu)$ is given by
@@
Other results
In this section, we provide more simulation results in the main text. Specifically, Tables 7, 8 and 9 show the simulation results of Section 3.2 of the main text for the scenarios:
  • $\mathbf{c:d=50:50}$ and one-sided test (less).
  • $\mathbf{c:d=50:50}$ and one-sided test (greater).
  • $\mathbf{c:d=20:80}$ and two-sided test.
Table 10 show the simulation results of Section 3.2 of the main text for the normal assumption. Figures 10, 11 and 12 show the results of Section 3.3 in main text for 3 combinations of $(\alpha,\nu)$: $(1,5)$, $(2,5)$, and $(3,5)$, respectively. Finally, Table 11 lists the index of brain regions in our real data analysis.